The pH of buffer solution prepared by mixing 50 ml of 0.2 molar acetic acid and 25 ml of Sodium Acetate is 4.8 then the concentration of Sodium Acetate is pka Of Acetic Acid equals to 4.8

Grade:11

1 Answers

Arun
5 years ago
Concentrations in the mixture: 

[CH3COOH] = 0.050*0.2/ 0.075 = 0.133 M 
[CH3COO-] = 0.025*x/0.075 = x/3 

pH = pKa + log[CH3COO-]/[CH3COOH] 

pKa = 4.76 
pH = 4.8 

0.04 = log[(x/3)/0.133] 
1.10 = x/0.40 
x = 0.44 

[CH3COO-] = x/3 = 0.15 M